For the reaction

2CH4(g)⇌C2H2(g)+3H2(g)2CH4(g)⇌C2H2(g)+3H2(g)

KcKcK_c = 0.170 at 1760. ∘C∘C . What is KpKpK_p for the reaction at this temperature?
Answer #1

Consider reaction, 2 CH4 (g) C2H2 (g) +3 H2 (g)

For above reaction , K c = 0.170 at 1760 0 C.

We have relation, K p = K c ( R T ) n

n = ( 1 + 3 ) - 2 = 4 -2 = 2

Substituting T = 1760 + 273 = 2033 K , R = 0.082057 L atm mol -1 K -1 , K c = 0.170 and n = 2 in above relation,we get

K p = ( 0.170 ) ( 0.082057 2033 ) 2

K p = 4731

ANSWER : K p = 4.73 10 3